# Refining the WASP-132 multi-planetary system: discovery of a cold giant planet and mass measurement of a hot super-Earth

### v1

@@ 4.1 Comparison to previous studies
We find our stellar parameters in agreement with Hord et al. (2022) and other previous studies. Hord et al. (2022) found a radius of 10.05 0.28 and mass of 121.9 21.9 for WASP-132 b, while Hellier et al. (2017) found a mass of 130.3 9.5 . Notably Hord et al. (2022) included a fit for WASP-132 c in their RV analysis to obtain the mass of WASP-132 b. With a measurement 136.1 4.8 , we decrease by a factor of two the mass uncertainty for WASP-132 b, and we find a similar radius of 10.10 0.43 . Hord et al. (2022) found a radius of 1.85 0.10 for WASP-132 c, which is similar to our result of 1.841 0.094 .
@@ 4.2 Dynamical history
Although WASP-132 c and WASP-132 b are not in resonance, we performed a Transit Timing Variation (TTV) analysis of WASP-132 b to test if any other non-transiting or non-RV detected planets were causing perturbations of the planet. We conducted an analysis of the TTVs for WASP-132 b with juliet by utilizing all available photometric datasets from TESS. Instead of fitting a single period () and time-of-transit center (), juliet employed a method that seeks individual transit times. This approach involved fitting each transit independently and determining one transit time for each one, resulting in a more consistent and coherent analysis. Figure 8 illustrates the results of our analysis, which compares the observed transit times with the calculated linear ephemeris derived from all the transits. Notably, no significant variations were detected within the data.
@@
The stellar obliquity or the observed excess of spin-orbit misalignment can put important constraints on the formation pathways of hot Jupiters. If WASP-132b did migrate early-on in the proto-planetary disk, measuring the spin-orbit angles of the two planets using the Rossiter-McLaughlin effect will allow the co-planarity of their orbits and their possible misalignments with the star to be assessed which constrains their dynamical history. As noted previously, ESPRESSO spectroscopic observations of the transits of both planets were obtained and are planned for publication by a separate team.
@@ 4.3 Astrometric signal
Astrometry can be an ideal method to provide exact masses of companions in combination with RVs. A commonly applied approach analyses the proper motion anomaly between Hipparcos (ESA, 1997) and Gaia (Gaia Collaboration et al., 2016b), described in more details by Snellen & Brown (2018) and Brandt et al. (2021). WASP-132 was too faint to be observed by Hipparcos, which means that the comparison is not possible; however, DR3 (Gaia Collaboration et al., 2023) includes an additional 12 months of observations more than DR2 (Gaia Collaboration et al., 2018b) and can be compared. The short-period planets are averaged into the proper motions reported by these two datasets, but the proper motion anomaly of WASP-132 d (1817 days) should be visible, if significant. Whereas the long term trend we detected in the RVs should be seen as a constant proper motion offset and will likely not introduce any difference between DR2 and DR3.
@@
There is indeed a substantial difference (mu = 0.30 mas yr-1) between the proper motion values in Right Ascension from DR3 (mu = 12.26 0.02 mas yr-1) and DR2 (mu = 12.56 0.07 mas yr-1). While the difference in proper motions in Declination (muδ = 0.04 mas yr-1) is of the same order as the error margins (DR3 muδ = 73.17 0.02 mas yr-1 and DR2 muδ = 73.21 0.07 mas yr-1). As the proper motion of DR3 also includes the data of DR2, the difference in proper motion is an indication of astrometric signal. However, drawing conclusions about the exact mass requires extensive modeling beyond the scope of this paper.
Additionally, Gaia DR3 reports a Renormalized Unit Weight Error (RUWE) of 1.23, where 1 is generally seen as a good measurement and above 1.4 as binarity. A RUWE above 1.2, however, can already indicate a proper motion anomaly. In other words, Gaia detects the astrometric signal of a companion but does not necessarily confirm it as a binary system. The excess noise level, at 0.11, while relatively low, ideally should be zero, indicating modeling errors in either the excess noise associated with the source or the attitude noise. This excess noise is reported with a high significance (20.9). In conclusion, given the companion’s minimal mass of 5.16 MJup and the presence of a significant astrometric signal, it is plausible that WASP-132 d may be a brown dwarf.
4.4 Internal characterization of the two giant planets
@@
The available measurements of the mass and the radius of the planet WASP-132 b make it possible to infer its bulk heavy-element content with giant planet evolution models (e.g., Miller & Fortney, 2011; Thorngren et al., 2016; Müller & Helled, 2023a). This is an important piece of information since it can be compared to predictions from formation models and therefore be used as an additional constraint (e.g., Guillot et al., 2006; Johansen & Lambrechts, 2017; Hasegawa et al., 2018). With a K, WASP-132 b is significantly cooler than strongly irradiated hot Jupiters. This makes the inferred composition less uncertain since the planet should not be inflated (e.g., Fortney et al., 2021). Here, we used the giant planet evolution models from Müller & Helled (2021) to calculate the cooling of the planet. Fig. 9 shows the radius as a function of time assuming that the planet has zero metallicity. Comparing it to the observed radius, it is clear that the planet is not inflated compared to predictions from the model, and that it must be enriched with heavy elements.
@@
To quantify the enrichment, we inferred the planet’s bulk heavy-element mass fraction in a Monte-Carlo fashion (see e.g., Müller & Helled, 2023b, for a review): we create sample planets by drawing from the observed planetary mass, radius, and age distributions. For each sample, we calculate the evolution for a range of bulk metallicities in order to find which value would match the sampled radius at the right age. By repeating this process, the posterior distribution of the planet’s heavy-element mass can be estimated. The result is shown in Fig. 10. The posterior is roughly Gaussian, with a mean of and a standard deviation of . This yields a heavy-element mass of , which is similar in magnitude to what would be expected from the critical core mass in core-accretion models (e.g., Helled et al., 2014). For comparison, if the planet had stellar metallicity, its heavy-element mass would be .
Since WASP-132 d does not have a measured radius, it is not possible to infer its bulk composition in the same way. However, we can provide a range of plausible radii by calculating its evolution from the measured for a range of bulk metallicities. We limited the models to since the upper bound would mean that the planet has about , which would be very metal-rich. While such extreme enrichments exist, they are not very common (Thorngren et al., 2016). The cooling tracks are shown in Fig. 11.
@@ 4.5 Internal characterization of WASP-132 c
Figure 12 clearly shows that WASP-132 b sits slightly above the Earth-like composition line, suggesting a refractory-rich composition. To probe the scope of interior structures that are consistent with the observed constraints, we use the four layer model of Dorn et al. (2017). It includes an iron core, a rocky mantle, a water layer, and a hydrogen-helium atmosphere. Using a nested sampling algorithm (Buchner et al., 2014), we vary the layer masses, the elemental ratios of Mg/Si and Fe/Si in the mantle and the age of the planet, taking the stellar values for the priors. In order to put upper bounds on the water and atmospheric mass fractions we ran a model assuming no atmosphere (no-atmosphere model) and a model assuming no water layer (no-water model), respectively. Additionally, we ran a model where we put no constraints on the layer masses (free model). Our results are listed in Table 5.
@@
All nested sampling models favors a refractory-rich composition which is dominated by metals and silicates and a negligible atmosphere. It should be noted, however, that the assumed internal structure is very simple. In reality, the internal structure of planets can be more complex, including and enrichment of the atmosphere with heavy elements and the hydration of the mantle and/or core. Accounting for more sophisticated structure could affect the inferred results. However, our main conclusions that WASP-132c is an exoplanet dominated by refractory material are unlikely to change.
@@ 4.6 WASP-132: a rare multi-planet system
Figure 13 shows a Mass-Radius diagram highlighting multi-planetary systems with precise densities ( 25 and 8) from the PlanetS catalog (Otegi et al. 2020), and in particular systems composed of a hot or warm Jupiter (P100 dyas) and a nearby small planet (R 4 ). Among the 127 planets in 80 systems, only 6 have this particular configuration: WASP-132, WASP-47 (Bryant & Bayliss 2022), 55 Cnc (Bourrier et al. 2018), KOI-94 (Weiss et al. 2013), WASP-84 (Maciejewski et al. 2023) and KOI-142 (Weiss et al. 2020). This once again demonstrates the rare contribution of this system, and the mass measurement of WASP-132 c, to a very small existing population.
The small planets with accompanying hot Jupiters consist of five super-Earths and two sub-Neptunes. Considering the sub-Neptunes, WASP-47 d is located at a longer period than its hot Jupiter companion, while the super-Earth WASP-47 e is inside the hot Jupiter. KOI-142 b orbits with a much longer period than the other planets in this diagram (P 10.92 days), explaining its larger size due to lower irradiation. Its associated Jupiter is also less ”hot”.
The super-Earths with hot or warm Jupiters are on the more massive end of this planet type and are likely composed of refractory elements. This may reveal that we are biased to detect the heaviest super-Earths in the presence of a hot Jupiter and lighter inner rocky planets in this kind of system may exist. Smaller planets like Kepler-730 c (R = 1.57 0.13 ; Cañas et al. 2019), which have no mass measurement, could be lighter. The lack of a similar system between the the sub-Neptunes and super-Earths with hot Jupiters could be intimately linked to their short orbital period and therefore to the irradiation received, stripping these planets of their atmosphere. Hot Jupiters then play an important role behind, limiting an internal planet’s period.
@@ 5 Conclusion
We report refined bulk measurements of the WASP-132 system. The 7.1-day hot Jupiter WASP-132 b was first discovered by WASP (Hellier et al., 2017) and an inner 1.01-day super-Earth was subsequently discovered by TESS (Hord et al., 2022). Here we also report the discovery of an 2.7 AU outer giant planet and long-term linear trend in the RVs suggestive of another outer companion. Using over nine years of CORALIE observations and highly-sampled HARPS data we determine the masses of the planets from smallest to largest orbital period to be Mp = , Mp = , and M = , respectively. Using TESS and CHEOPS photometry data we measure the radii of the two inner transiting planets to be and . We also performed an independent stellar analysis to find M⋆ = and R⋆ = . We find that the hot Jupiter WASP-132 b is not inflated and is enriched with heavy elements. For the super-Earth WASP-132 c, our internal structure modeling favors a refractory-rich composition that is dominated by metals and silicates and a negligible atmosphere. We find that both the inner 1-day super-Earth and hot Jupiter are likely in circular orbits with eccentricities of 0. Given the presence of a nearby inner-planet and its low-eccentricity, the hot Jupiter WASP-132 b likely migrated through quieter mechanisms than high-eccentricity migration making it a rare contribution to the current planet population.
@@ Acknowledgements.
We thank the Swiss National Science Foundation (SNSF) and the Geneva University for their continuous support to our planet low-mass companion search programs. This work was carried out in the frame of the Swiss National Centre for Competence in Research (NCCR) supported by the SNSF including grants 51NF40_182901 and 51NF40_205606. This work used computations that were performed at the University of Geneva on the ”Yggdrasil” High Performance Computing (HPC) clusters. This publication makes use of The Data & Analysis Center for Exoplanets (DACE), which is a facility based at the University of Geneva (CH) dedicated to extrasolar planet data visualization, exchange, and analysis. DACE is a platform of NCCR and is available at https://dace.unige.ch. This paper includes data collected by the TESS mission. Funding for the TESS mission is provided by the NASA Explorer Program. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. V.A. acknowledges the support by FCT - Fundação para a Ciência e a Tecnologia through national funds (grants: UIDB/04434/2020, UIDP/04434/2020, and 2022.06962.PTDC). DJA is supported by UKRI through the STFC (ST/R00384X/1) and EPSRC (EP/X027562/1). NCS acknowledges the funding by the European Union (ERC, FIERCE, 101052347). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work was supported by FCT - Fundação para a Ciência e a Tecnologia through national funds and by FEDER through COMPETE2020 - Programa Operacional Competitividade e Internacionalização by these grants: UIDB/04434/2020; UIDP/04434/2020.References
- ESA (1997) 1997, ESA Special Publication, Vol. 1200, The HIPPARCOS and TYCHO catalogues. Astrometric and photometric star catalogues derived from the ESA HIPPARCOS Space Astrometry Mission
@@
- Gaia Collaboration et al. (2018a) Gaia Collaboration, Brown, A. G. A., Vallenari, A., et al. 2018a, A&A, 616, A1
- Gaia Collaboration et al. (2018b) Gaia Collaboration, Brown, A. G. A., Vallenari, A., et al. 2018b, A&A, 616, A1
@@
- Gaia Collaboration et al. (2016a) Gaia Collaboration, Prusti, T., de Bruijne, J. H. J., et al. 2016a, A&A, 595, A1
- Gaia Collaboration et al. (2016b) Gaia Collaboration, Prusti, T., de Bruijne, J. H. J., et al. 2016b, A&A, 595, A1
@@
- Sha et al. (2022) Sha, L., Vanderburg, A. M., Huang, C. X., et al. 2022, arXiv e-prints, arXiv:2209.14396
@@
- Sousa (2014) Sousa, S. G. 2014, [arXiv:1407.5817] [arXiv:1407.5817]
@@
- Winn (2010) Winn, J. N. 2010, arXiv e-prints, arXiv:1001.2010
@@
- Zhu (2023) Zhu, W. 2023, arXiv e-prints, arXiv:2306.16691

### v2

@@ 4.1 Comparison to previous studies
We find our stellar parameters in agreement with Hord et al. (2022) and other previous studies. Hord et al. (2022) found a radius of 10.05 0.28 and mass of 121.9 21.9 for WASP-132 b, while Hellier et al. (2017) found a radius of 9.75 0.34 and mass of 130.3 9.5 . Notably Hord et al. (2022) included a fit for WASP-132 c in their RV analysis to obtain the mass of WASP-132 b. With a measurement 136.1 4.8 , we decrease by a factor of two the mass uncertainty for WASP-132 b, and we find a similar radius of 10.10 0.43 . Hord et al. (2022) found a radius of 1.85 0.10 for WASP-132 c, which is similar to our result of 1.841 0.094 . For the two inner planets we find bulk densities of = g cm-3 for WASP-132 c and = g cm-3 for WASP-132 b.
@@ 4.2 Dynamical history
Although WASP-132 c and WASP-132 b are not in resonance, we performed a TTV analysis of WASP-132 b to test if any other non-transiting or non-RV detected planets were causing perturbations of the planet. We conducted an analysis of the TTVs for WASP-132 b with juliet by utilizing all available photometric datasets from TESS. Instead of fitting a single period () and time-of-transit center (), juliet employed a method that seeks individual transit times. This approach involved fitting each transit independently and determining one transit time for each one, resulting in a more consistent and coherent analysis. We also included the TRAPPIST observation of the WASP-132 b transit made on 2014 May 5 from Hellier et al. (2017) for our TTV analysis. Figure 8 illustrates the results of our analysis, which compares the observed transit times with the calculated linear ephemeris derived from all the transits. Notably, no significant variations were detected within the TESS data and the TRAPPIST transit is only a 1 outlier.
@@
The stellar obliquity or the observed excess of spin-orbit misalignment can put important constraints on the formation pathways of hot Jupiters. If WASP-132b did migrate early on in the proto-planetary disk, measuring the spin-orbit angles of the two planets using the Rossiter-McLaughlin effect will allow the coplanarity of their orbits and their possible misalignments with the star to be assessed, which constrains their dynamical history. As noted previously, ESPRESSO spectroscopic observations of the transits of both planets were obtained and are planned for publication by a separate team.
@@ 4.3 Astrometric signal
Astrometry can be an ideal method to provide exact masses of companions in combination with RVs. A commonly applied approach analyses the proper motion anomaly between Hipparcos (ESA, 1997) and Gaia (Gaia Collaboration et al., 2016), described in more details by Snellen & Brown (2018) and Brandt et al. (2021). WASP-132 was too faint to be observed by Hipparcos, which means that the comparison is not possible; however, DR3 (Gaia Collaboration et al., 2023) includes an additional 12 months of observations more than DR2 (Gaia Collaboration et al., 2018) and can be compared. The short-period planets are averaged into the proper motions reported by these two datasets, but the proper motion anomaly of WASP-132 d (1817 days) should be visible, if significant. Whereas the long term trend we detected in the RVs should be seen as a constant proper motion offset and will likely not introduce any difference between DR2 and DR3.
@@
There is a 4-sigma difference (mu = 0.30 0.072 mas yr-1) between the proper motion values in Right Ascension from DR3 (mu = 12.26 0.02 mas yr-1) and DR2 (mu = 12.56 0.07 mas yr-1). While the difference in proper motions in Declination (muδ = 0.04 0.072 mas yr-1) is not significant (DR3 muδ = 73.17 0.02 mas yr-1 and DR2 muδ = 73.21 0.07 mas yr-1). As the proper motion of DR3 also includes the data of DR2, the difference in proper motion could be an indication of astrometric signal but it could also be caused by improved Gaia calibration and data reduction.
Additionally, Gaia DR3 reports a Renormalized Unit Weight Error (RUWE) of 1.23, where 1 is generally seen as a good measurement and above 1.4 as a possible indication of binarity (Kervella et al., 2022). The Gaia DR3 astrometric excess noise level of 0.11, while relatively low, ideally should be zero. Astrometric excess noise can indicate modeling errors in either the excess noise associated with the source or the attitude noise. These hints of astrometric variations must be interpreted with caution and drawing conclusions is beyond the scope of this paper. Gaia DR4 will enable the determination of WASP-132 d’s true mass and its relative orbital inclination with respect to the two inner transiting planets. We note that WASP-132 d would need to have an orbital inclination larger than 66 degrees from an edge-on orbit to have a mass greater than 13 .
4.4 Internal characterization of the giant planet WASP-132 b
@@
The available measurements of the mass and the radius of the planet WASP-132 b make it possible to infer its bulk heavy-element content with giant planet evolution models (e.g., Miller & Fortney, 2011; Thorngren et al., 2016; Müller & Helled, 2023a). This is an important piece of information since it can be compared to predictions from formation models and therefore be used as an additional constraint (e.g., Guillot et al., 2006; Johansen & Lambrechts, 2017; Hasegawa et al., 2018). With a K, WASP-132 b is significantly cooler than strongly irradiated hot Jupiters. This makes the inferred composition less uncertain since the planet should not be inflated (e.g., Fortney et al., 2021). Here, we used the giant planet evolution models from Müller & Helled (2021) to calculate the planet’s cooling. Figure 9 shows the radius as a function of time assuming that the planet has a bulk metallicity between zero and and 25%. Comparing it to the observed radius, it is clear that the planet is not inflated compared to predictions from the model, and that it must be enriched with heavy elements.
@@
To quantify the enrichment, we inferred the planet’s bulk heavy-element mass fraction in a Monte-Carlo fashion (see e.g., Müller & Helled, 2023b, for a review): we create sample planets by drawing from the observed planetary mass, radius, and age distributions. For each sample, we calculate the evolution for a range of bulk metallicities to find which value would match the sampled radius at the right age. By repeating this process, the posterior distribution of the planet’s heavy-element mass can be estimated. The result is shown in Fig. 10. The posterior is roughly Gaussian, with a mean of and a standard deviation of . This yields a heavy-element mass of , which is similar in magnitude to what would be expected from the critical core mass in core-accretion models (e.g., Helled et al., 2014). For comparison, if the planet had stellar metallicity, its heavy-element mass would be .
@@ 4.5 Internal characterization of WASP-132 c
Figure 11 shows a Mass-Radius diagram highlighting multi-planetary systems with precise densities ( 25 and 8) from the PlanetS catalog666https://dace.unige.ch/exoplanets/ (Otegi et al. 2020; Parc et al. 2024), and in particular systems composed of a hot or warm Jupiter (P100 days) and a nearby small planet (R 4 ). Figure 11 clearly shows that WASP-132 c sits slightly above the Earth-like composition line, suggesting a refractory-rich composition. To probe the scope of interior structures that are consistent with the observed constraints, we use the four layer model of Dorn et al. (2017). It includes an iron core, a rocky mantle, a water layer, and a hydrogen-helium atmosphere. Using a nested sampling algorithm (Buchner et al., 2014), we vary the layer masses, the elemental ratios of Mg/Si and Fe/Si in the mantle and the age of the planet, taking the stellar values for the priors. In order to put upper bounds on the water and atmospheric mass fractions we ran a model assuming no atmosphere (no-atmosphere model) and a model assuming no water layer (no-water model), respectively. Additionally, we ran a model where we put no constraints on the layer masses (free model). Our results are listed in Table 5.
@@
All nested sampling models favor a refractory-rich composition dominated by metals and silicates and a negligible atmosphere. It should be noted, however, that the assumed internal structure is very simple. In reality, the internal structure of planets can be more complex, including and enrichment of the atmosphere with heavy elements and the hydration of the mantle and/or core. Accounting for more sophisticated structure could affect the inferred results. However, our main conclusions that WASP-132c is an exoplanet dominated by refractory material are unlikely to change.
@@ 4.6 WASP-132: A rare multi-planet system
As noted, Fig. 11 shows a mass-radius diagram highlighting multi-planetary systems with precise densities from the PlanetS catalog. Among the 127 planets in 80 systems, only six have a hot or warm Jupiter (P100 days) and a nearby small planet (R 4 ): WASP-132, WASP-47 (Bryant & Bayliss 2022), 55 Cnc (Bourrier et al. 2018), KOI-94 (Weiss et al. 2013), WASP-84 (Maciejewski et al. 2023) and KOI-142 (Weiss et al. 2020). This once again demonstrates the rare contribution of this system, and the mass measurement of WASP-132 c, to a very small existing population.
The small planets with accompanying hot Jupiters consist of five super-Earths and two sub-Neptunes. Considering the sub-Neptunes, WASP-47 d is located at a longer period than its hot Jupiter companion, while the super-Earth WASP-47 e is inside the hot Jupiter. KOI-142 b orbits with a much longer period than the other planets in this diagram (P 10.92 days), explaining its larger size due to lower irradiation. Its associated Jupiter is also less ”hot.”
The super-Earths with hot or warm Jupiters are on the more massive end of this planet type and are likely composed of refractory elements. This may reveal that we are biased to detect the heaviest super-Earths in the presence of a hot Jupiter and lighter inner rocky planets in this kind of system may exist. Smaller planets like Kepler-730 c (R = 1.57 0.13 ; Cañas et al. 2019), which have no mass measurement, could be lighter. The lack of a similar system between the super-Earths and the sub-Neptunes could be intimately linked to their short orbital period and therefore to the irradiation received, stripping these planets of their atmosphere. Hot Jupiters then play an important role behind, limiting an internal planet’s period.
@@ 5 Conclusion
We report refined bulk measurements of the WASP-132 system. The 7.1-day hot Jupiter WASP-132 b was first discovered by WASP (Hellier et al., 2017) and an inner 1.01-day super-Earth was subsequently discovered by TESS (Hord et al., 2022). Here we also report the discovery of an 2.7 AU outer giant planet and long-term linear trend in the RVs suggestive of another outer companion. Using over nine years of CORALIE observations and highly sampled HARPS data, we determined the masses of the planets from smallest to largest orbital period to be Mp = , Mp = , and M = , respectively. Using TESS and CHEOPS photometry data, we measured the radii of the two inner transiting planets to be and . We also performed an independent stellar analysis to find M⋆ = and R⋆ = . We find that the hot Jupiter WASP-132 b is not inflated but enriched with heavy elements. For the super-Earth WASP-132 c, our internal structure modeling favors a refractory-rich composition that is dominated by metals and silicates and a negligible atmosphere. We find that both the inner 1-day super-Earth and hot Jupiter are likely in circular orbits with eccentricities of 0. Given the presence of a nearby inner planet and its low eccentricity, the hot Jupiter WASP-132 b likely migrated through quieter mechanisms than high-eccentricity migration, making it a rare contribution to the current planet population.
6 Data availability
A table of the reduced CHEOPS photometric observations is available in a machine-readable format at The Strasbourg astronomical Data Center (CDS) via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/. The reduced TESS photometric observations are available in a machine-readable format at the CDS. The CORALIE RVs and all activity indicator data are available in a machine-readable format at the CDS. The HARPS RVs and all activity indicator data are available in a machine-readable format at the CDS. Figures mentioned in the text as supplementary material are published on Zenodo at https://doi.org/10.5281/zenodo.14271139.
@@ Acknowledgements.
We thank the Swiss National Science Foundation (SNSF) and the Geneva University for their continuous support to our planet low-mass companion search programs. This work was carried out in the frame of the Swiss National Centre for Competence in Research (NCCR) supported by the SNSF including grants 51NF40_182901 and 51NF40_205606. This work used computations that were performed at the University of Geneva on the ”Yggdrasil” High Performance Computing (HPC) clusters. This publication makes use of The Data & Analysis Center for Exoplanets (DACE), which is a facility based at the University of Geneva (CH) dedicated to extrasolar planet data visualization, exchange, and analysis. DACE is a platform of NCCR and is available at https://dace.unige.ch. This paper includes data collected by the TESS mission. Funding for the TESS mission is provided by the NASA Explorer Program. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. V.A. acknowledges the support by FCT - Fundação para a Ciência e a Tecnologia through national funds (grants: UIDB/04434/2020, UIDP/04434/2020, and 2022.06962.PTDC). DJA is supported by UKRI through the STFC (ST/R00384X/1) and EPSRC (EP/X027562/1). NCS acknowledges the funding by the European Union (ERC, FIERCE, 101052347). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work was supported by FCT - Fundação para a Ciência e a Tecnologia through national funds and by FEDER through COMPETE2020 - Programa Operacional Competitividade e Internacionalização by these grants: UIDB/04434/2020; UIDP/04434/2020. N.G. thanks Daniel Thomas Watson for his pertinacious and adamantine support.References
@@
- ESA (1997) ESA, ed. 1997, ESA Special Publication, Vol. 1200, The HIPPARCOS and TYCHO catalogues. Astrometric and photometric star catalogues derived from the ESA HIPPARCOS Space Astrometry Mission
- Espinoza (2018) Espinoza, N. 2018, Research Notes of the American Astronomical Society, 2, 209
@@
- Gaia Collaboration et al. (2018) Gaia Collaboration, Brown, A. G. A., Vallenari, A., et al. 2018, A&A, 616, A1
@@
- Gaia Collaboration et al. (2016) Gaia Collaboration, Prusti, T., de Bruijne, J. H. J., et al. 2016, A&A, 595, A1
@@
- Gomes da Silva et al. (2011) Gomes da Silva, J., Santos, N. C., Bonfils, X., et al. 2011, A&A, 534, A30
@@
- Parc et al. (2024) Parc, L., Bouchy, F., Venturini, J., Dorn, C., & Helled, R. 2024, A&A, 688, A59
@@
- Sha et al. (2023) Sha, L., Vanderburg, A. M., Huang, C. X., et al. 2023, MNRAS, 524, 1113
@@
- Sousa (2014) Sousa, S. G. 2014, in Determination of Atmospheric Parameters of B, ed. E. Niemczura, B. Smalley, & W. Pych, 297–310
@@
- Wilson (1978) Wilson, O. C. 1978, ApJ, 226, 379
- Winn (2010) Winn, J. N. 2010, in Exoplanets, ed. S. Seager, 55–77
@@
- Zhu (2024) Zhu, W. 2024, Research in Astronomy and Astrophysics, 24, 045013